Accountancy

Suresh a  partner in a firm withdrawns Rs 6000 regularly on the last day of every month. Interest on 12% premium. Calculate interest on drawings 
Answer #1

Interest on drawings = 6000×12/100=72000

72000×5.5/12×12/100=3960

answered by: Advika